what is g-6=2?
can you show me the steps?

Answer by elima(1433)

g-6=2
The first step would be to get the g by itself, so we will + 6 to each side.
We are adding 6 because we want to do the opposite opperation to make it 0.
The operation is subtract, the opposite of subtract is add.
g -6+6=2+6
And remember whatever we do to one side we have to do to the other.
answer:
g=8
Hope this helps =)